# A variational discretization method for mean curvature flows by the Onsager principle

### v1

@@ Abstract
The mean curvature flow describes the evolution of a surface(a curve) with normal velocity proportional to the local mean curvature. It has many applications in mathematics, science and engineering. In this paper, we develop a numerical method for mean curvature flows by using the Onsager principle as an approximation tool. We first show that the mean curvature flow can be derived naturally from the Onsager variational principle. Then we consider a piecewisely linear approximation of the curve and derive a discrete geometric flow. The discrete flow is described by a system of ordinary differential equations for the nodes of the discrete curve. We prove that the discrete system preserve the energy dissipation structure in the framework of the Onsager principle and this implies the energy decreasing property. The ODE system can be solved by an improved Euler scheme and this leads to an efficient fully discrete scheme. We first consider the method for a simple mean curvature flow and then extend it to volume preserving mean curvature flow and also a wetting problem on substrates. Numerical examples show that the method has optimal convergence rate and works well for all the three problems.
@@ 1 Introduction
[…]
@@
In this paper, we develop a new finite element method for mean curvature flow by using the Onsager principle as an approximation tool. We consider three different problems, namely the standard mean curvature flow, the volume conserving mean curvature flow and a wetting problem with contact with substrate. We show that the dynamic equations of all the three problems can be derived naturally by the Onsager principle. Furthermore, by considering piecewisely linear approximation of the surface, we can derive semi-discrete problems which preserve the energy dissipation relations just as the continuous problems. The discrete volume can also be preserved if the continuous problem preserve the volume. The semi-discrete problems can be further discretized in time by an improved Euler scheme. Numerical examples show that the method works well for all the three problems. In particular, it has optimal convergence rate in space for a mean curvature flow with analytic solutions.
@@
For simplicity in presentation, we consider only two dimensional problems in the paper. However, the method can be generalized to higher dimensional problems straightforwardly. The structure of the paper is as follows. In Section 2, we introduce the main idea to use the Onsager variational principle as an approximation tool. In Section 3, we first derive a continuous partial differential equation for a simple smooth closed curve under mean curvature flow using the Onsager principle. Subsequently, we discretize the equation using piecewisely linear curves and prove the discrete energy decays property. We introduce a penalty term to the energy functional to avoid mesh distortion. In Section 4, we further consider a mean curvature flow with the constraint of volume conservation. We derive a continuity equation and also a discrete problem. In Section 5, we further generalize the method to solve a wetting problem, which is a non-closed curve contacting with substrates. In Section 6, we present several numerical examples to demonstrate the efficiency of the methods. We give a few conclusion remarks in the final section.
@@ 2 The Onsager principle as an approximation tool
$\Phi(\dot{u})=\frac{\xi}{2}\left\|{\dot{u}}\right\|_{2}^{2},$ (2)
@@
˙$\dot{\mathcal{E}}(u;\dot{u})=\langle\frac{\delta\mathcal{E}(u)}{\delta u},\dot{u}\rangle=\left.\frac{\text{d}\mathcal{E}(u+\lambda\dot{u})}{\text{d}\lambda}\right|_{\lambda=0}.$ (3)
@@
[…]
@@
$(\xi\dot{u},\varphi)=-\langle\frac{\delta\mathcal{E}(u)}{\delta u},\varphi\rangle,\quad\forall\varphi\in V,$ (6)
@@
[…]
Recently, the Onsager principle is used to derive an approximate model for a physical system [Doi2015]. The key idea is follows. Suppose that the system can be described by a system of slow variables $\mathbf{a}=(a_{1},a_{2},\cdots,a_{N})$. Suppose we can compute the approximate energy $\mathcal{E}_{h}(\mathbf{a})$ and the approximate dissipation function $\Phi_{h}(\dot{\mathbf{a}})$. Then we can derive a discrete ODE system by the Onsager principle, which is given by
@@
The method has been used to develop reduced models in complicated two-phase flows [XuXianmin2016, DiYana2016], some soft matter problems [ManXingkun2016, ZhouJiajia2017], solid dewetting [Jiang19b], and also in dislocation dynamics [dai2021boundary]. It can also be used to derive numerical method for wetting problems [lu2021efficient]. In this following, we will investigate the application of the Onsager principle in deriving numerical schemes for mean curvature flows.
@@ 3.1 Derivation of the continuous equation by the Onsager principle
For simplicity in presentation, we consider only the two dimensional case. For a simple smooth closed curve $\Gamma(t)$ in the plane, the points on it are denoted by $\vec{x}(t)$. We aim to derive the mean curvature flow equation of $\Gamma(t)$ by using the Onsager variational principle.
@@
[…]
$\frac{\text{d}}{\text{d}t}\int_{\Gamma(t)}f\text{d}s=\int_{\Gamma(t)}[\frac{\partial f}{\partial t}+\nabla_{\Gamma}\cdot(f\vec{v})]\text{d}s=\int_{\Gamma(t)}(\frac{\text{d}f}{\text{d}t}+f\cdot\kappa v_{n})\text{d}s.$ (9)
@@
$\Phi=\frac{1}{2}\int_{\Gamma(t)}v_{n}^{2}\text{d}s.$ (11)
@@
$\mathcal{R}=\Phi+\dot{\mathcal{E}}=\int_{\Gamma(t)}(\frac{1}{2}v_{n}^{2}+\kappa v_{n})\text{d}s.$ (12)
Using the Onsager principle, we minimize the Rayleighian $\mathcal{R}$ with respect to $v_{n}$, i.e.
$\min_{v_{n}}\mathcal{R}.$ (13)
@@
$v_{n}+\kappa=0.$ (14)
Substituting the equation (14) into (10), we have
˙$\dot{\mathcal{E}}=\int_{\Gamma(t)}\kappa v_{n}\text{d}s=-\int_{\Gamma(t)}\kappa^{2}\text{d}s=-2\Phi(v_{n})\leq 0,$ (15)
@@ 3.2 A discrete mean curvature flow by the Onsager principle
In this section, we aim to numerically solve the dynamic equation (14). Suppose that the curve $\Gamma(t)$ can be approximated by piecewise linear segments, as illustrated in Figure 1. Denote the nodes by
$\vec{x}_{1}(t),\vec{x}_{2}(t),...,\vec{x}_{n}(t),$ (17)
where $\vec{x}_{i}(t)=(x_{i}^{(1)}(t),x_{i}^{(2)}(t))^{T}\ (1\leq i\leq n)$. The line segments connecting adjacent nodes are denoted by $\Gamma_{i}(t)=\overline{\vec{x}_{i}(t)\vec{x}_{i+1}(t)}\ (1\leq i\leq n-1)$ and $\Gamma_{n}(t)=\overline{\vec{x}_{n}(t)\vec{x}_{1}(t)}$. Therefore, the approximate curve is given by $\Gamma_{h}(t)=\bigcup\limits_{i=1}^{n}\Gamma_{i}(t)$.
Figure 1: Selection of interpolation points
@@
$\vec{x}_{h}(t)$ $=\mu[\vec{x}_{i+1}(t)-\vec{x}_{i}(t)]+\vec{x}_{i}(t)=(1-\mu)\vec{x}_{i}(t)+\mu\vec{x}_{i+1}(t),$ (18)
where $0\leq\mu=\frac{s}{|\vec{x}_{i+1}-\vec{x}_{i}|}\leq 1$ is a parameter. The velocity of each point $\vec{x}_{h}(t)$ is given by
$\vec{v}_{h}(t)=(1-\mu)\dot{\vec{x}}_{i}(t)+\mu\dot{\vec{x}}_{i+1}(t).$ (19)
@@
$\mathcal{E}_{h}=\int_{\Gamma_{h}(t)}\text{d}s=\sum_{i=1}^{n}\int_{\Gamma_{i}(t)}\text{d}s=\sum_{i=1}^{n}|\Gamma_{i}(t)|=\sum_{i=1}^{n}|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|.$ (22)
@@
$\Phi_{h}=\frac{1}{2}\int_{\Gamma_{h}(t)}v_{h,n}^{2}\text{d}s=\frac{1}{2}\sum_{i=1}^{n}\int_{\Gamma_{i}(t)}[\vec{v}_{h}(t)\cdot\vec{n}_{i}]^{2}\text{d}s.$ (23)
By the Onsager principle, we minimize the discrete Rayleighian function with respect to $\dot{\vec{x}}_{i}$, i.e.
@@
We can specify the explicit formula for (25) by direct calculations. Denote by $\dot{X}=(\dot{\vec{x}}_{1}^{T},\dot{\vec{x}}_{2}^{T},...,\dot{\vec{x}}_{n}^{T})^{T}$ and notice that $\Phi_{h}$ is a quadratic positive definite function with respect to $\dot{\vec{x}}_{i}\ (1\leq i\leq n)$. The equation set (25) can be rewritten as
@@
where the coefficient matrix $A$ is symmetric positive definite and the right-hand term $G=(\vec{g}_{1}^{T},\vec{g}_{2}^{T},...,\vec{g}_{n}^{T})^{T}$, such that $\Phi_{h}=\frac{1}{2}\dot{X}^{T}A\dot{X}$ and $\vec{g}_{i}=-\frac{\partial\mathcal{E}_{h}}{\partial\vec{x}_{i}}$. The explicit formulas for $A$ and $G$ are given in Appendix.
@@
Suppose $\dot{X}=(\dot{\vec{x}}_{1}^{T},\dot{\vec{x}}_{2}^{T},...,\dot{\vec{x}}_{n}^{T})^{T}$ is the solution of the system (26), then we have
$\frac{\text{d}\mathcal{E}_{h}}{\text{d}t}=-2\Phi_{h}(\dot{X})\leq 0,$ (27)
@@
$\frac{\text{d}\mathcal{E}_{h}}{\text{d}t}=\sum_{i=1}^{n}\frac{\partial\mathcal{E}_{h}}{\partial\vec{x}_{i}}\cdot\dot{\vec{x}}_{i}=-\sum_{i=1}^{n}\vec{g}_{i}\cdot\dot{\vec{x}}_{i}=-\dot{X}^{T}G=-\dot{X}^{T}A\dot{X}=-2\Phi_{h}(\dot{X}).$ (28)
Here we have used the equation (26). Since the coefficient matrix $A$ is positive definite, we prove the theorem.
@@ 3.3 Stabilization for uniform distribution of vertexes
$\mathcal{E}_{h}^{\delta}$ $=|\Gamma_{h}(t)|+\delta\sum_{i=1}^{n}\bigl{(}\frac{|\Gamma_{i}(t)|}{|\Gamma_{i+1}(t)|}-1\bigr{)}^{2}$ (29)
$=\sum_{i=1}^{n}[|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|+\delta\bigl{(}\frac{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}{|\vec{x}_{i+2}(t)-\vec{x}_{i+1}(t)|}-1\bigr{)}^{2}].$
@@
$\frac{\partial\Phi_{h}}{\partial\dot{\vec{x}}_{i}}+\frac{\partial\mathcal{E}_{h}^{\delta}}{\partial\vec{x}_{i}}=0.$ (31)
Similarly the equation (31) can be rewritten as
@@
where $G^{\delta}=((\vec{g}_{1}^{\delta})^{T},(\vec{g}_{2}^{\delta})^{T},...,(\vec{g}_{n}^{\delta})^{T})^{T}$ with $\vec{g}_{i}^{\delta}=-\frac{\partial\mathcal{E}_{h}^{\delta}}{\partial\vec{x}_{i}}$. We refer to the Appendix for the explicit forms of $g_{i}^{\delta}$.

### v3

@@ Abstract
The mean curvature flow describes the evolution of a surface (a curve) with normal velocity proportional to the local mean curvature. It has many applications in mathematics, science and engineering. In this paper, we develop a numerical method for mean curvature flows by using the Onsager principle as an approximation tool. We first show that the mean curvature flow can be derived naturally from the Onsager variational principle. Then we consider a piecewise linear approximation of the curve and derive a discrete geometric flow. The discrete flow is described by a system of ordinary differential equations for the nodes of the discrete curve. We prove that the discrete system preserve the energy dissipation structure in the framework of the Onsager principle and this implies the energy decreasing property. The ODE system can be solved by the improved Euler scheme and this leads to an efficient fully discrete scheme. We first consider the method for a simple mean curvature flow and then extend it to the volume preserving mean curvature flow and also a wetting problem on substrates. Numerical examples show that the method has optimal convergence rate and works well for all the three problems.
@@ 1 Introduction
where $\vec{v}$ denotes the velocity of motion of the surface, and $H$ denotes the mean curvature of the surface. $\vec{n}$ is the unit normal vector of the surface. The mean curvature flow first appeared in materials science in the 1920s, and the general mathematical model was first proposed in 1957 in [40], where Mullins used it to describe how grooves develop on the surfaces of hot polycrystals. The mean curvature flow has found numerous applications. One is in image smoothing, where it is used to improve image quality by removing unnecessary noise and enhancing specific features without altering the transmitted information [26]. Additionally, the mean curvature flow has been employed to model a special form of the reaction-diffusion equation in chemical reactions [44].
The numerical approximation of mean curvature flow originates from the pioneering work of Dziuk in 1990 [19]. He proposed a parametric finite element method (PFEM) to solve the problem based on the observation that the mean curvature flow is a diffusion equation for the surface embedding in a bulk region. However, as the surface evolves, the nodes in the PFEMs may overlap, leading to mesh distortion [27]. Subsequently, various techniques have been introduced in PFEMs that allow points on interpolated curves or surfaces to shift tangentially to produce a better distribution of nodes. For example, Deckelnick and Dziuk introduced an artificial tangential velocity in their work [14]. By similar motivation, Elliott and Fritz employed a DeTurck’s trick to introduce a tangential velocity [37]. A widely used method is the parametric finite element method developed by Barrett, Garcke, and Nürnberg [5, 6, 7]. They utilize a different variational form which can induce tangential velocity for the mesh nodes automatically. The method have been generalized some other geometrical flows [8, 9, 10, 28, 4, 35, 2, 3, 27]. Recently, there are significant progresses in rigorous convergence analysis for discrete schemes for mean curvature flows and some other geometric flows [30, 31, 32, 33, 1]. In addition, there exist many other methods to approximate the mean curvature flow numerically, such as the threshold dynamics method [39, 45, 22, 50, 48], the level set method [43, 24, 12, 47, 52, 46], and the phase field method [20, 23, 21, 25], etc.
@@
In this paper, we develop a new finite element method for mean curvature flow by using the Onsager principle as an approximation tool. We consider three different problems, namely the standard mean curvature flow, the volume preserving mean curvature flow and a wetting problem with contact with substrate. We show that the dynamic equations of all the three problems can be derived naturally by the Onsager principle. Furthermore, by considering a piecewise linear approximation of the surface, we can derive semi-discrete problems which preserve the energy dissipation relations just as the continuous problems. The discrete volume can also be preserved if the continuous problem preserve the volume. The semi-discrete problems can be further discretized in time by the improved Euler scheme. Numerical examples show that the method works well for all the three problems. In particular, it has optimal convergence rate in space for a mean curvature flow with analytic solutions. We mainly consider mean curvature flows of a curve in two dimensional space, also known as curve shortening flow in literature. Our method can be generalized to higher dimensional problems while there will be more restrictions on the time discrete schemes to avoid mesh distortion.
@@
The structure of the paper is as follows. In Section 2, we introduce the main idea to use the Onsager variational principle as an approximation tool. In Section 3, we first derive a continuous partial differential equation for a simple smooth closed curve under mean curvature flow using the Onsager principle. Subsequently, we discretize the equation using piecewise linear curves and prove the discrete energy decays property. We introduce a penalty term to the energy functional to avoid mesh distortion. We also briefly discuss the extension of the method to higher dimensional cases. In Section 4, we further consider a mean curvature flow with the constraint of volume conservation. We derive a continuity equation and also a discrete problem. In Section 5, we further generalize the method to solve a wetting problem, which is a non-closed curve contacting with substrates. In Section 6, we present several numerical examples to demonstrate the efficiency of the methods. We give a few conclusion remarks in the final section.
@@ 2 The Onsager principle as an approximation tool
$\Phi(\dot{u})=\frac{\xi}{2}\left\|{\dot{u}}\right\|_{2}^{2},$
@@
˙$\dot{\mathcal{E}}(u;\dot{u})=\left\langle\frac{\delta\mathcal{E}(u)}{\delta u},\dot{u}\right\rangle=\left.\frac{\text{d}\mathcal{E}(u+\lambda\dot{u})}{\text{d}\lambda}\right|_{\lambda=0}.$
@@
Based on the above aforementioned definitions, the Onsager variational principle can be stated as follows [41, 42, 17]. For any dissipative system without inertial effect, the evolution of the system can be obtained by minimizing the Rayleighian with respect to $\dot{u}$. In other words, the evolution equation of $u$ is obtained by
@@
$(\xi\dot{u},\varphi)=-\left\langle\frac{\delta\mathcal{E}(u)}{\delta u},\varphi\right\rangle,\quad\forall\varphi\in V,$ (4)
@@
In physics, Eq. (5) implies a balance between the frictional force $-\xi\dot{u}$ and the driven force $\frac{\delta\mathcal{E}(u)}{\delta u}$. For a more comprehensive introduction of the Onsager principle, we refer to [17].
Recently, the Onsager principle is used to derive an approximate model for a physical system [18]. The key idea is follows. Suppose that the system can be described by a system of slow variables $\mathbf{a}=(a_{1},a_{2},\cdots,a_{N})$. Suppose we can compute the approximate energy $\mathcal{E}_{h}(\mathbf{a})$ and the approximate dissipation function $\Phi_{h}(\dot{\mathbf{a}})$. Then we can derive a discrete ODE system by the Onsager principle, which is given by
@@
The method has been used to develop reduced models in complicated two-phase flows [49, 16], some soft matter problems [38, 54], solid dewetting [29], and also in dislocation dynamics [13]. It can also be used to derive numerical method for wetting problems [36]. In this following, we will investigate the application of the Onsager principle in deriving numerical schemes for mean curvature flows.
@@ 3.1 Derivation of the continuous equation by the Onsager principle
We mainly consider the mean curvature flow in two dimensional space. For a simple smooth closed curve $\Gamma(t)$ in the plane, the points on it are denoted by $\vec{x}(t)$. We aim to derive the mean curvature flow equation of $\Gamma(t)$ by using the Onsager variational principle.
@@
where $s$ is the arc length parameter. In order to calculate the changing rate of the surface energy with respect to time, we need to employ the following transport equation for a closed curve [11]
$\frac{\text{d}}{\text{d}t}\int_{\Gamma(t)}f\text{d}s=\int_{\Gamma(t)}\left(\frac{\partial f}{\partial t}+v_{n}\vec{n}\cdot\nabla f+f\kappa v_{n}\right)\text{d}s.$ (7)
@@
$\Phi=\frac{1}{2}\int_{\Gamma(t)}|\vec{v}|^{2}\text{d}s.$
@@
$\mathcal{R}=\Phi+\dot{\mathcal{E}}=\int_{\Gamma(t)}\left(\frac{1}{2}|\vec{v}|^{2}+\kappa v_{n}\right)\text{d}s.$
Using the Onsager principle, we minimize the Rayleighian $\mathcal{R}$ with respect to $\vec{v}$, i.e.
$\min_{\vec{v}}\mathcal{R}.$
@@
$\vec{v}=-\kappa\vec{n}.$ (9)
This is the mean curvature flow [34]. Substituting Eq. (9) into (8), we have
˙$\dot{\mathcal{E}}=\int_{\Gamma(t)}\kappa v_{n}\text{d}s=-\int_{\Gamma(t)}\kappa^{2}\text{d}s=-2\Phi\leq 0,$
@@ 3.2 A discrete mean curvature flow by the Onsager principle
In this section, we aim to numerically solve the dynamic equation (9). Suppose that the curve $\Gamma(t)$ can be approximated by piecewise linear segments, as illustrated in Fig. 1. Denote the nodes by
$\vec{x}_{1}(t),\vec{x}_{2}(t),...,\vec{x}_{n}(t),$
where $\vec{x}_{i}(t)=(x_{i}^{(1)}(t),x_{i}^{(2)}(t))^{\top}\ (1\leq i\leq n)$. The line segments connecting adjacent nodes are denoted by $\Gamma_{i}(t)=\overline{\vec{x}_{i}(t)\vec{x}_{i+1}(t)}\ (1\leq i\leq n-1)$ and $\Gamma_{n}(t)=\overline{\vec{x}_{n}(t)\vec{x}_{1}(t)}$. Therefore, the approximate curve is given by $\Gamma_{h}(t)=\bigcup\limits_{i=1}^{n}\Gamma_{i}(t)$.
Figure 1: Selection of interpolation points.
@@
$\vec{x}_{h}(t)$ $=\mu[\vec{x}_{i+1}(t)-\vec{x}_{i}(t)]+\vec{x}_{i}(t)=(1-\mu)\vec{x}_{i}(t)+\mu\vec{x}_{i+1}(t),$
where $0\leq\mu=\frac{s}{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}\leq 1$ is a parameter and $s$ is the arc length parameter of $\Gamma_{i}(t)$. $\vec{x}_{n+1}$ is set to be $\vec{x}_{1}$. The velocity of each point $\vec{x}_{h}(t)$ is given by
$\vec{v}_{h}(t)=(1-\mu)\dot{\vec{x}}_{i}(t)+\mu\dot{\vec{x}}_{i+1}(t).$
@@
$\mathcal{E}_{h}=\int_{\Gamma_{h}(t)}\text{d}s=\sum_{i=1}^{n}\int_{\Gamma_{i}(t)}\text{d}s=\sum_{i=1}^{n}|\Gamma_{i}(t)|=\sum_{i=1}^{n}|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|.$
@@
$\Phi_{h}=\frac{1}{2}\int_{\Gamma_{h}(t)}|\vec{v}_{h}|^{2}\text{d}s.$
By the Onsager principle, we minimize the discrete Rayleighian with respect to $\dot{\vec{x}}_{i}$, i.e.
@@
We can specify the explicit formula for (11) by direct calculations. Denote by $\dot{X}=(\dot{\vec{x}}_{1}^{\top},\dot{\vec{x}}_{2}^{\top},...,\dot{\vec{x}}_{n}^{\top})^{\top}$ and notice that $\Phi_{h}$ is a quadratic function with respect to $\dot{\vec{x}}_{i}\ (1\leq i\leq n)$. The equation set (11) can be rewritten as
@@
where the coefficient matrix $A$ is symmetric positive definite and the right-hand term $G=(\vec{g}_{1}^{\top},\vec{g}_{2}^{\top},...,\vec{g}_{n}^{\top})^{\top}$, such that $\Phi_{h}=\frac{1}{2}\dot{X}^{\top}A\dot{X}$ and $\vec{g}_{i}=-\frac{\partial\mathcal{E}_{h}}{\partial\vec{x}_{i}}$. The explicit formulas for $A$ and $G$ are given in Appendix.
@@
Suppose $\dot{X}=(\dot{\vec{x}}_{1}^{\top},\dot{\vec{x}}_{2}^{\top},...,\dot{\vec{x}}_{n}^{\top})^{\top}$ is the solution of the system (12), then we have
$\frac{\text{d}\mathcal{E}_{h}}{\text{d}t}=-2\Phi_{h}\leq 0,$
@@
$\frac{\text{d}\mathcal{E}_{h}}{\text{d}t}=\sum_{i=1}^{n}\frac{\partial\mathcal{E}_{h}}{\partial\vec{x}_{i}}\cdot\dot{\vec{x}}_{i}=-\sum_{i=1}^{n}\vec{g}_{i}\cdot\dot{\vec{x}}_{i}=-\dot{X}^{\top}G=-\dot{X}^{\top}A\dot{X}=-2\Phi_{h}.$
Here we have used Eq. (12). Since the coefficient matrix $A$ is positive definite, we prove the theorem.
@@ 3.3 Stabilization for uniform distribution of vertexes
$\mathcal{E}_{h}^{\delta}$ $=|\Gamma_{h}(t)|+\delta\sum_{i=1}^{n}\left(\frac{|\Gamma_{i}(t)|}{|\Gamma_{i+1}(t)|}-1\right)^{2}$ (13)
$=\sum_{i=1}^{n}\left[|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|+\delta\left(\frac{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}{|\vec{x}_{i+2}(t)-\vec{x}_{i+1}(t)|}-1\right)^{2}\right].$
@@
$\frac{\partial\Phi_{h}}{\partial\dot{\vec{x}}_{i}}+\frac{\partial\mathcal{E}_{h}^{\delta}}{\partial\vec{x}_{i}}=0.$
Similarly, the above equation can be rewritten as
@@
where $G^{\delta}=((\vec{g}_{1}^{\delta})^{\top},(\vec{g}_{2}^{\delta})^{\top},...,(\vec{g}_{n}^{\delta})^{\top})^{\top}$ with $\vec{g}_{i}^{\delta}=-\frac{\partial\mathcal{E}_{h}^{\delta}}{\partial\vec{x}_{i}}$. We refer to the Appendix for the explicit forms of $g_{i}^{\delta}$.
3.4 Mean curvature flow in high dimensional spaces
In the end of the section, we briefly discuss the mean curvature flow in higher dimensional spaces. For a n-dimensional hypersurface $\Gamma(t)$ in $\mathbb{R}^{n+1}$, we can also derive the mean curvature flow equation and its discretization by the Onsager variational principle.
$\mathcal{E}=\int_{\Gamma(t)}\text{d}s=|\Gamma(t)|.$
Similarly to the curve case, we compute the changing rate of the total energy by using the Reynolds transport equation as
˙$\dot{\mathcal{E}}=\frac{\text{d}}{\text{d}t}\int_{\Gamma(t)}\text{d}s=\int_{\Gamma(t)}Hv_{n}\text{d}s,$ (16)
where $v_{n}=\vec{v}\cdot\vec{n}$ is the normal velocity and $H=\sum_{j=1}^{n}\kappa_{j}$ is the mean curvature of $\Gamma(t)$ with $\kappa_{1},...,\kappa_{n}$ being the principal curvatures  [^1] Notice that the definition of the mean curvature differs from the common one $H=\frac{1}{n}\sum_{j=1}^{n}\kappa_{j}$ [15]. . Suppose the dissipation function in the system is given by
$\Phi=\frac{1}{2}\int_{\Gamma(t)}|\vec{v}|^{2}\text{d}s.$
Then the corresponding Rayleighian is defined as
$\mathcal{R}=\Phi+\dot{\mathcal{E}}=\int_{\Gamma(t)}\left(\frac{1}{2}|\vec{v}|^{2}+Hv_{n}\right)\text{d}s.$
By the Onsager principle, we minimize the Rayleighian $\mathcal{R}$ with respect to $\vec{v}$, i.e.
$\min_{\vec{v}}\mathcal{R}.$
The corresponding Euler-Lagrange equation is
$\vec{v}=-H\vec{n}.$ (17)
This gives a mean curvature flow equation in high dimensional space.
To discretize Eq. (17), we can follow the same approach as the curve case. We first approximate the surface $\Gamma(t)$ by a discrete surface $\Gamma_{h}(t):=\bigcup_{T\in\mathcal{T}_{h}}T$ consists of piecewise nondegenerate n-dimensional simplices $T$. Then we derive a discrete system for the vertexes of the triangulation by using the Onsager principle similarly as that in the previous subsections. In principle, we could solve the discrete system to approximate the mean curvature flow for the hypersurface. However, there may be more restrictions on the time discrete schemes to avoid mesh distortion in high dimensional cases. Some properly designed semi-implicit schemes may have nice property as the parametric finite element method [5]. This will be left for future work.